Let E be the solid bounded by y = x^2, z = 0, y + 2z = 4. Express the integral


1
Expert's answer
2021-12-01T11:07:39-0500
"\int\int\int_Ef(x,y,z)dV"

"=\displaystyle\int_{-2}^{2}\displaystyle\int_{0}^{x^2}\displaystyle\int_{0}^{2-y/2}f(x,y,z)dzdydx"